If α=α1i+α2j+α3keta=\ eta1i+\ eta2j+\ eta3k, thenαdot \etaαdot\ eta is ____?

a.\\(\\alpha_{1}\\beta_{1}+\\alpha_{2}\\beta_{2}+\\alpha_{3}\\beta_{3}\\)
b.\\(\\alpha_{1}\\beta_{1-\\alpha}\\beta_{2-\\alpha}\\beta_{3-\\alpha}\\)
c.\\(\\alpha_{1}\\beta_{1}-\\alpha_{2}\\beta_{2}-\\alpha_{3}\\beta_{3}\\)
d.\\(\\alpha_{1}\\beta_{1}+\\alpha_{2}\\beta_{2}-\\alpha_{3}\\beta_{3}\\)
1
Expert's answer
2020-03-23T11:54:18-0400

We are given "\overrightarrow{\alpha}=\alpha_1i+\alpha_2j+\alpha_3k" and "\overrightarrow{\beta}=\beta_1i+\beta_2j+\beta_3k"

Then the dot product of both vector is "\overrightarrow{\alpha}.\overrightarrow{\beta}"

"\overrightarrow{\alpha}.\overrightarrow{\beta}=\alpha_1\beta_1+\alpha_2\beta_2+\alpha_3\beta_3"